Sun 6:11 PM a 66% E- uni The charges and coordinates of two charged partides held fixed in an xy plane are q1 3.02 C, x1 3.56 cm, yi-0.970 cm and and (b) direction (with respect to +x-axis in the range (-180°;1801) of the electrostatic force on should a third particle of charge 93-3.53 uC be placed such that the net electrostatic force on particle 2 2 due to particle 1. At what (c) x and (d) y coordinates due to particles 1 and 3 is zero? (a) Number (b) Number (c) Number (d) Number
Answer #1

-2.62 cm, 2.48 cm 21 (3.56 cm, 0.97 cm) 0

Distance from particle 2 due to particle 1 which will be given as -

r21 = sqrt{}(x2 - x1)2 + (y2 - y1)2

r21 = sqrt{}[(-0.0262 m) - (0.0356 m)]2 + [(0.0248 m) - (0.0097 m)]2

r21 = sqrt{}0.00404725 m2

r21 = 0.063 m

(a) Magnitude of an electrostatic force on particle 2 due to particle 1 which will be given by -

| F21 | = ke |q1| |q2| / r212

| F21 | = [(9 x 109 Nm2/C2) (3.02 x 10-6 C) (6.20 x 10-6 C)] / (0.00404725 m2)

| F21 | = 41.6 N

(b) Direction of an electrostatic force on particle 2 due to particle 1 which will be given by -

tan heta = [(y2 - y1) / (x2 - x1)]

tan heta = { [(0.0248 m) - (0.0097 m)] / [(-0.0262 m) - (0.0356 m)] }

heta = tan-1 (-0.2443)

heta = - 13.7 degree

{ below the positive x-axis }

We know that, | F23 | = | F21 |

ke |q2| |q3| / r232 = ke |q1| |q2| / r212

r232 = r212 |q3| / |q1|

r23 = sqrt{}[(0.00404725 m) (3.53 x 10-6 C)] / (3.02 x 10-6 C)

r23 = sqrt{}0.00473 m2

r23 = 0.0687 m

converting m into cm :

r23 = 6.87 cm

(c) At what x coordinates should a third particle of charge q3 be placed?

x3 = x2 - r23 cos heta

x3 = (-2.62 cm) - (6.87 cm) cos 13.70

x3 = [(-2.62 cm) - (6.67 cm)]

x3 = - 9.29 cm

(d) At what y coordinates should a third particle of charge q3 be placed?

y3 = y2 - r23 sin heta

y3 = (2.48 cm) - (6.87 cm) sin 13.70

y3 = [(2.48 cm) + (1.62 cm)]

y3 = 4.10 cm